OpenStudy (anonymous):

A boy is twice as tall as his little sister and 30cm shorter than his father. Their combined height is 3.8m. Find the height of each person.

OpenStudy (anonymous):

Boy height = x sister height = x/2 father height = x+30 So: \[x + \frac{x}{2} + x + .30 = 3.8\]

OpenStudy (anonymous):

Then you can solve for x

OpenStudy (anonymous):

Then find the other heights

OpenStudy (anonymous):

THANK YOU